OpenStudy (anonymous):

Given: f(x) = 3 - x; g(x) = -2x Find f[g(2)].

OpenStudy (tyteen4a03):

Plug the numbers in: g(2) = -2(2) = -4. Now solve for f(x) with the result of g(2).

OpenStudy (anonymous):

f[g(2)] = f(-4) = 3 - (-4) =7